OpenStudy (anonymous):

Find the mean absolute deviation of the set of data 16, 20, 24, 25, 28, 29, 31 and 35.

OpenStudy (anonymous):

Add them together for the mean. it equals 209, then divide 209 by how many numbers there are which is 8 so that equals 26.125. that is your mean. no clue what absolute deciation means. sorry!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Mean Absolute Deviation = Sum ( x_i - mean) / n

OpenStudy (anonymous):

mean absolute deviation is the mean absolute distance from the mean. for each data value do:\[|x-x_m|\]where \[x_m = mean = 26.125\] then find the average of these as a formula: \[M.A.D = \frac{\sum_{i =1}^{i=n}|x_i - x_m|}{n}\]

OpenStudy (anonymous):

@Chlorophyll it has to be absolute value of (x_i - mean) i think

OpenStudy (anonymous):

Find the mean absolute deviation of the set of data 16, 20, 24, 25, 28, 29, 31 and 35. 1) 4.65 2) 4.70 3) 4.75 4) 4.80

OpenStudy (anonymous):

Yes, it should have absolute sign!

OpenStudy (anonymous):

i needdddd answer please

OpenStudy (anonymous):

do you understand our explanation?

OpenStudy (anonymous):

Why don't you input the numbers : start with |16 - 26.125 | + | 20 - 26.125 | + ....

OpenStudy (anonymous):

@tjdenerson It's not as bad as you imagine, just enter 8 subtractions with plus signs, then divide by 8! That's it :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Did you get the result?

OpenStudy (anonymous):

) 4.75 4) 4.80

OpenStudy (anonymous):

4.75

OpenStudy (anonymous):

tha

OpenStudy (anonymous):

Is it s..oo..o ha..rd.. :P